# Beyond the Encoder: Joint Encoder-Decoder Contrastive Pre-Training Improves Dense Prediction

### v2

@@ 4.2 Ablation studies
Table 5 presents an ablation study of the components in our DeCon-ML-L framework. The results show that while pre-training with a decoder and deep-supervision provides a modest benefit to downstream fine-tuning performance on COCO, even when only the encoder is transferred, the main performance improvement arises from the use of channel dropout in conjunction with decoder deep supervision. This combination yields the most substantial gains, highlighting the critical role of channel dropout when paired with deep supervision in our approach.
Table 5: Ablation of the different components of DeCon-ML-L and their impact on the downstream performance. We averaged all fine-tuning results over three runs. $\Delta$ is the performance gap between the previous framework and the framework updated with the corresponding new component.
@@
Table 6 shows DeCon-ML performance with different values of encoder loss weight $\alpha$ and channel dropout. A channel dropout probability of 0.5 provided the best downstream performance for DeCon-ML-L. Hence, this value was chosen to ablate $\alpha$. An $\alpha$ value of 0 gave the best downstream performance meaning that the decoder loss can replace the encoder one and pre-train efficiently an encoder on its own. For DeCon-SL, an $\alpha$ value of 0.25 resulted in the best downstream performance, suggesting that having an encoder loss at the bottleneck is still beneficial to the pre-training in the absence of skip connections.
Table 6: Impact of the channel dropout probability and of the encoder loss weight $\alpha$ on DeCon. We pre-trained DeCon adaptations of SlotCon framework on COCO dataset using different parameters and evaluate the pre-trained encoder fine-tuning on downstream tasks that use the same decoder architecture. We averaged all fine-tuning results over three runs.
Abla tion Frame work Deco der $\alpha$ Drop. Obj. Det. Inst. Seg. Frame work Deco der $\alpha$ Sem. Seg.
@@
Dropout
DeCon-ML-L
FPN
0.50 0 40.85 36.83 - - - - -
0.25 40.95 36.97 - - - - -
0.50 41.10 37.05 - - - - -
0.75 41.02 37.05 - - - - -
$\alpha$
0.75 0.50 40.98 36.90
DeCon-SL (No Drop)
FCN
0.75 71.98 76.18
0.50 41.10 37.05 0.50 72.42 75.62
0.25 41.05 37.13 0.25 73.01 76.21
0 41.18 37.11 0 72.17 75.71
@@ 5 Discussion
Table 7: COCO and COCO+ pretraining downstream performance improvement with DeCon-ML-L and Cohen’s d value. PixCon performance was extracted from the original paper.
Method COCO obj det. COCO Inst. Seg. VOC Sem. Seg. City Sem. Seg.
$\Delta$AP $\Delta$AP $\Delta$mIoU $\Delta$mIoU
COCO Pre-training Performance Improvement $\Delta$ (calculated from the respective papers)
SlotCon (2022) (Comp. to prev. works) +0.50 +0.40 -0.40 +0.10
PixCon (2024) (Comp. to SlotCon) 0.00 +0.06 +1.30 +0.51
DeCon-ML-L (Comp. to SlotCon) +0.37 +0.32 +1.42 +0.50
Cohen’s d (DeCon-ML-L VS SlotCon) 2.35 1.94 6.37 2.51
COCO+ Pre-training Performance Improvement $\Delta$ (calculated from the respective papers)
SlotCon (2022) (Comp. to prev. works) +0.60 +0.50 Not avail. Not avail.
PixCon (2024) (Comp. to SlotCon) -0.50 -0.50 -0.20 +0.40
DeCon-ML-L (Comp. to SlotCon) +0.45 +0.27 +1.43 +0.57
Cohen’s d (DeCon-ML-L VS SlotCon) 5.42 3.21 5.25 1.22
@@
In this paper, we introduced a contrastive framework adaptation by incorporating a decoder contrastive loss to an existing framework. We showed that our proposed SSL adaptation outperforms previous state-of-the-art methods on COCO object detection and instance segmentation and achieves comparable results on Pascal VOC semantic segmentation when pre-trained on COCO dataset, and consistently outperforms previous methods when pre-trained on COCO+, and ImageNet except for Cityscapes semantic segmentation (Tab. 1). This adaptation enhances the representation of the pre-trained encoder, improving its transfer performance for dense downstream tasks, regardless of the homogeneity between pre-training and fine-tuning decoders (as shown in Tab. 1).
We introduced a combined loss where encoder and decoder losses contribute to the total loss in a weighted manner.
Figure 2: Pre-training loss dynamics for DeCon-ML-L and SlotCon.
@@
Figure 2 illustrates the loss behavior across different pre-trainings when both the encoder and decoder are used. As shown, decoder pre-training does not influence the encoder’s loss dynamics, supporting the fact that the encoder and decoder loss terms are conceptually non-competing. We also observed that increasing the weight for the decoder loss compared to that of the encoder enhances the performance of the pre-trained models on downstream tasks. This could be attributed to the non-competing nature of an encoder and decoder loss compared to approaches that utilize a reconstruction loss for a decoder and a different loss for the encoder in an encoder-decoder architecture [28]. Additionally, Fig. 3 presents the slots learned by SlotCon encoder and by the different decoder levels of DeCon-ML. It demonstrates that the decoder learns features that capture similar concepts as the encoder’s, while being more spatially precise.
@@
Figure 3: Slots, as defined in SlotCon, learned at different outputs of our architecture. The top left image from COCO validation dataset is overlayed with the slot that was the most represented in original SlotCon’s encoder output feature map, resized to the image’s shape. Other slots displayed are the slots that had the biggest overlap with SlotCon’s main encoder output slot. More details on the slots creation are available in Supplementary Material 6.
@@
Moreover, we noticed that the encoder loss is optional when pre-training with DeCon-ML, demonstrating the efficiency of the decoder deep supervision and channel dropout combination to ensure a comprehensive pre-training of both encoder and decoder. The lack of skip-connections in DeCon-SL with FCN decoder makes the combined encoder-decoder loss preferable to ensure the best pre-training. In Supplementary Material Section 7, we provide experiments on using DeCon-SL with an FPN decoder. We observed that with an FPN decoder, dropout does not provide the same benefits in DeCon-SL compared to DeCon-ML. Hence, channel dropout should only be applied to the different encoder levels when pre-training with a skip-connected decoder and decoder deep supervision (DeCon-ML). Framework specific hyperparameter tuning of the decoder auxiliary layers could also increase the fine-tuning performance (shown in Supplementary Material Section 8).
@@
We demonstrated that our method is effective with a different type of encoders. Specifically, pre-training a ConvNeXt-small backbone, which is both more efficient and larger than the ResNet-50 backbone, also resulted in superior downstream performance when integrated with a decoder in the DeCon-ML-L framework. These results attest to the scalability and practical utility of the proposed approach for pre-training modern backbone architectures.
@@
DeCon was also used to successfully adapt three different SSL frameworks using different concepts (pixel level VS object/group level) to learn local information (Tab. 3). This suggests that the proposed DeCon adaptation could be used in other SSL contrastive methods designed for dense downstream tasks.
@@
Table 8: Pre-training parameter count, GPU setup, and time required per epoch for the different architectures. We study various pre-training setups on the COCO dataset using the SlotCon framework. Module sizes are given in millions of parameters.
Pre-training Enc. size Dec. size Enc. Aux. Layers Size Dec. Aux. Layers Size Full Archi. Size GPU setup Time per epoch (s)
SlotCon 23.51 x 2 0 71.51 0 118.52 1 * A100 80GB 255
DeCon-SL-FCN ($\alpha$ = 0.25) 23.51 x 2 5.31 x 2 71.51 56.83 185.97 1 * A100 80GB 300
DeCon-ML-S ($\alpha$ = 0) 23.51 x 2 3.35 x 2 0 59.08 110.03 1 * A100 80GB 280
DeCon-ML-L ($\alpha$ = 0) 23.51 x 2 3.35 x 2 0 227.31 281.02 8 * H100 80GB 165
@@
DeCon adaptation of an SSL framework provided notable advantages for fine-tuning on small-scale datasets and out-of-domain dense prediction tasks, outperforming the encoder-only baseline (see Tab. 4). This improvement persisted even when pre-training was conducted directly on the small target datasets. Furthermore, DeCon achieved superior performance compared to related frameworks when only a limited fraction (5–25%) of the original labeled data was available for fine-tuning. These results highlight the value of DeCon in settings with very limited annotated samples or small-scale, domain-specific datasets, such as those in medical or agricultural applications, regardless of whether the pre-trained decoder is transferred.
@@
Performance of deep learning models can vary across multiple runs due to inherent stochasticity. To mitigate the impact of randomness while considering computational resource constraints, each experiment was conducted three times, with all fine-tunings for a specific downstream task performed on similar hardware. The performance improvements achieved with our DeCon framework compared to the original framework are statistically significant, as evidenced by large Cohen’s $d$ values (>0.8>0.8> 0.8) in Tab. 7, exceeding two standard deviations (see Supplementary Material Section 9), and $p$-values less than 0.05 (0.012 for COCO object detection and 0.048 for instance segmentation) from Wilcoxon signed-rank tests. Details on the computation of these values are provided in Supplementary Material 10. Furthermore, the performance gains shown in Tab. 7 are comparable to those reported in prior studies.
@@
One limitation of our approach is its high memory and computational demand. However, we have shown in Tab. 1 that if GPU is a constraint, the user can decide to reduce the number of decoder levels used in the pre-training and tune the decoder projector layers to reduce the parameter count of DeCon-ML architecture and match the parameter count of the original framework (as shown in Tab. 8), while still improving the performance of the pre-training.
@@
Brempong et al. [4] followed a staged approach, first pre-training the encoder with supervision on annotated datasets, then pre-training the decoder with denoising. However, this multi-stage process relies on labeled data. It could also result in suboptimal feature alignment and increased training complexity. Future works could involve extending our proposed DeCon adaptation to offer a multi-stage continual pre-training in a contrastive manner.
@@ 6 Conclusion
In this paper, we introduced DeCon—a novel contrastive framework adaptation that enhances the performance of self-supervised learning approaches for dense predictions tasks, such as object detection, semantic segmentation, and instance segmentation. We showed that this adaptation improves the encoder’s representation significantly. We proposed two variants of DeCon: DeCon-SL and DeCon-ML. The former introduces a decoder contrastive loss, and the latter extends DeCon-SL by introducing channel dropout and decoder deep supervision to maximize the encoder’s pre-training power. By integrating the proposed contrastive joint encoder-decoder pre-training strategy into an existing framework, we show that DeCon consistently outperforms or matches state-of-the-art methods across diverse dense prediction tasks, highlighting the effectiveness of this unified pre-training approach.
@@ Acknowledgments
Computing resources were provided by the University of British Columbia Cloud Innovation Centre, powered by Amazon Web Services https://cic.ubc.ca/. This research was also partly enabled by support provided by the Digital Research Alliance of Canada https://alliancecan.ca/, and the Research Computing Services group at the University of Calgary https://rcs.ucalgary.ca/RCS_Home_Page.

### v3

@@ 4.1 Experiments
Table 6: ResNet-50 backbone transfer performance of SlotCon, DeCon-SL (FCN decoder) and DeCon-ML (FPN decoder), pre-trained on COCO and fine-tuned on agriculture datasets with 10% and 100% labeled data.
Framework PlantDoc (Obj. Det.) Detecting Diseases (Obj. Det.) PlantSeg (Sem. Seg.)
AP AP AP AP mIoU mIoU
10% 100% 10% 100% 10% 100%
Random Init. 6.59 19.81 19.14 34.76 16.96 24.96
SlotCon 17.53 38.37 26.54 48.45 20.72 28.78
DeCon-SL ($\alpha=0.25$) 17.06 39.94 26.81 48.85 20.56 29.39
DeCon-ML-S ($\alpha=0$, drop=0.5) 17.82 39.40 26.82 48.53 21.00 29.16
DeCon-ML-L ($\alpha=0$, drop=0.5) 17.98 40.03 27.16 48.57 20.66 29.77
@@ 4.2 Ablation studies
Table 7 presents an ablation study of the components in our DeCon-ML-L framework. The results show that while pre-training with a decoder and deep-supervision provides a modest benefit to downstream fine-tuning performance on COCO, the main performance improvement arises from the use of channel dropout in conjunction with decoder deep supervision. This combination yields the most substantial gains, highlighting the critical role of channel dropout when paired with deep supervision in our approach.
Table 7: Ablation of the different components of DeCon-ML-L and their impact on the downstream performance. $\Delta$ is the performance gap between the previous framework and the framework updated with the corresponding new component.
@@
Table 8 shows DeCon-ML performance with different values of encoder loss weight $\alpha$, channel dropout, and number of decoder levels. A channel dropout probability of 0.5 provided the best downstream performance for DeCon-ML-L. Hence, this value was chosen to ablate $\alpha$. An $\alpha$ value of 0 gave the best downstream performance meaning that the decoder loss can replace the encoder one and pre-train efficiently an encoder on its own. We also ablated with number of decoder levels in DeCon-ML-L where 4 decoder levels yielded the best averaged performance across tasks. For DeCon-SL, an $\alpha$ value of 0.25 resulted in the best downstream performance, suggesting that having an encoder loss at the bottleneck is still beneficial to the pre-training in the absence of skip connections.
Table 8: Ablation of channel dropout, encoder loss weight ($\alpha$), and decoder levels in DeCon, pre-trained on COCO for 800 epochs and evaluated by fine-tuning on downstream tasks.
Abla tion Frame work Deco der $\alpha$ Drop. Obj. Det. Inst. Seg. Abla tion Frame work Deco der $\alpha$ Dec Levels Sem. Seg.
@@
Dropout DeCon-ML-L (4 decoder levels) FPN 0.50 0 40.85 36.83 Dec. Level DeCon-ML-L (Drop = 0.5) FPN 0 2 72.70 76.07
0.25 40.95 36.97 3 73.14 76.10
0.50 41.10 37.05 4 72.92 76.45
0.75 41.02 37.05 - - - -
$\alpha$ 0.75 0.50 40.98 36.90 $\alpha$ DeCon-SL (No Drop) FCN 0.75 - 71.98 76.18
0.50 41.10 37.05 0.50 - 72.42 75.62
0.25 41.05 37.13 0.25 - 73.01 76.21
0 41.18 37.11 0 - 72.17 75.71
@@ 5 Discussion
In this paper, we introduced a contrastive framework adaptation by incorporating a decoder contrastive loss to an existing framework. We showed that our proposed SSL adaptation achieved SOTA performance in object detection, instance and semantic segmentation tasks when pre-trained on COCO+ and ImageNet-1K datasets, and in most tasks when pre-trained on COCO (Tab. 1). This adaptation enhances the representation of the pre-trained encoder, improving its transfer performance for dense downstream tasks, even in tasks beyond object detection, instance and semantic segmentation, regardless of the homogeneity between pre-training and fine-tuning decoders.
Figure 2: DeCon-ML-L and SlotCon pre-training loss dynamics.
@@
We introduced a combined loss where encoder and decoder losses contribute to the total loss in a weighted manner. Figure 2 illustrates losses behavior across different pre-trainings when both the encoder and decoder are used. As shown, decoder pre-training does not affect the encoder loss dynamics, supporting the fact that the encoder and decoder loss terms are conceptually non-competing. We also observed that increasing the weight for the decoder loss compared to that of the encoder enhances the performance of the pre-trained models. This could be attributed to the non-competing nature of the encoder and decoder losses compared to approaches that use a reconstruction loss for a decoder and a different loss for the encoder in an encoder-decoder architecture [31]. Additionally, Fig. 3 presents the slots learned by SlotCon encoder and by the different decoder levels of DeCon-ML. It demonstrates that the decoder learns features that capture similar concepts as the encoder’s, while being more spatially precise.
@@
Figure 3: Slots, as defined in SlotCon [55], learned at different outputs of our architecture. The top left image from COCO validation dataset is overlayed with the slot that was the most represented in original SlotCon’s encoder output feature map, resized to the image’s shape. Other slots displayed are the slots that had the biggest overlap with SlotCon’s main encoder output slot. More details on the slots creation are available in Supplementary Material 6.
@@
Moreover, we noticed that the encoder loss is optional when pre-training with DeCon-ML, demonstrating the efficiency of the decoder deep supervision and channel dropout combination to ensure a comprehensive pre-training of both encoder and decoder. The lack of skip-connections in DeCon-SL with FCN decoder makes the combined encoder-decoder loss preferable to ensure the best pre-training. In Supplementary Material Section 7, we provide experiments on using DeCon-SL with an FPN decoder. We observed that with an FPN decoder, dropout does not provide the same benefits in DeCon-SL compared to DeCon-ML. Hence, channel dropout should only be applied to the different encoder levels when pre-training with a skip-connected decoder and decoder deep supervision (DeCon-ML). Framework-specific hyperparameter tuning of the decoder auxiliary layers could also increase the fine-tuning performance (Supplementary Material Section 8).
@@
We demonstrated that our method is effective with different encoders. Specifically, pre-training a ConvNeXt-Small backbone, which is both more efficient and larger than the ResNet-50 backbone, also resulted in superior downstream performance when integrated with a decoder in the DeCon-ML-L framework. These results attest to the scalability and practical utility of the proposed approach for pre-training modern backbone architectures. Additionally, Tab. 3 shows that a ConvNeXt-S backbone pre-trained with DeCon-SL for dense tasks outperforms larger ViT-based SSL methods, despite using fewer pre-training epochs.
@@
DeCon was also used to successfully adapt three different SSL frameworks using different concepts (pixel level VS object/group level) to learn local information (Tab. 4). This suggests that the proposed DeCon adaptation could be used in other SSL contrastive methods targeting dense tasks.
@@
Table 9: Pre-training parameter count and computational cost for the different architectures. We study various pre-training setups on the COCO dataset using the SlotCon framework. Module sizes are given in millions of parameters. GPU-days are calculated on H100 80GB GPUs.
Pre-training Enc. size Dec. size Enc. Aux. Layers Size Dec. Aux. Layers Size Full Archi. Size GPU- days
SlotCon 23.51 x 2 0 71.51 0 118.52 1.53
DeCon-SL-FCN ($\alpha$ = 0.25) 23.51 x 2 5.31 x 2 71.51 56.83 185.97 1.77
DeCon-ML-S ($\alpha$ = 0) 23.51 x 2 3.35 x 2 0 59.08 110.03 1.78
DeCon-ML-L ($\alpha$ = 0) 23.51 x 2 3.35 x 2 0 227.31 281.02 11.00
@@
DeCon adaptation of SSL frameworks yields clear advantages on small-scale datasets and out-of-domain dense prediction tasks, outperforming encoder-only baselines (Tab. 5, Tab. 6). These gains persist even when pre-training on the small target datasets, or when fine-tuning with only a fraction of labeled data. This highlights DeCon’s value in domains with limited annotations, like medical and agricultural applications.
@@
Performance of deep learning models can vary across multiple runs due to inherent stochasticity. To mitigate the impact of randomness while considering computational resource constraints, each experiment was conducted three times, with all fine-tunings for a specific downstream task performed on similar hardware. The performance improvements achieved with our DeCon framework compared to the original framework are statistically significant, as evidenced by large Cohen’s $d$ values (>0.8>0.8) (in Supplementary Material Tab. S6), exceeding two standard deviations (see Supplementary Material Section 9), and $p$-values less than 0.05 (0.012 for COCO object detection and 0.048 for instance segmentation) from Wilcoxon signed-rank tests. Details on the computation of these values are provided in Supplementary Material Section 10. Furthermore, the performance gains, shown in Supplementary Material Tab. S6, are comparable to those reported in prior studies.
@@
A limitation of our approach is its high memory and computational demand. However, as shown in Tab. 1, when GPU access is limited, reducing the number of decoder levels and tuning the decoder projector layers can decrease the parameter count and training cost of DeCon-ML (e.g., DeCon-ML-S in Tab. 9) while still enhancing pre-training performance. It is also notable that DeCon-SL and DeCon-ML-S have comparable training costs as their baseline.
@@
Brempong et al. [5] followed a staged approach, first pre-training the encoder with supervision on annotated datasets, then pre-training the decoder with denoising. However, this multi-stage process relies on labeled data. It could also result in suboptimal feature alignment and increased training complexity. Future works could involve extending our proposed DeCon adaptation to offer a multi-stage continual pre-training in a contrastive manner. We also plan to adapt DeCon to ViT architectures in the future.
@@ 6 Conclusion
In this paper, we introduced DeCon—a novel contrastive framework adaptation that enhances the performance of SSL approaches for various dense predictions tasks. We showed that this adaptation improves the encoder’s representation significantly. We proposed two variants of DeCon: DeCon-SL and DeCon-ML. The former introduces a decoder contrastive loss, and the latter extends DeCon-SL by introducing channel dropout and decoder deep supervision to maximize the encoder’s pre-training power. By integrating the proposed contrastive joint encoder-decoder pre-training strategy into an existing framework, we show that DeCon consistently outperforms or matches state-of-the-art methods across diverse dense prediction tasks, highlighting the effectiveness of this unified pre-training approach.
@@ Acknowledgments
We extend our sincere gratitude to the University of Calgary VPR Office for their generous support through VPR Catalyst Grants, which was instrumental in making this research possible. We also acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC Discovery Grant RGPIN-2024-04966) and Alberta Innovates. This work was partly supported by Alberta Innovates Graduate Student Scholarship and T. Chen Fong Doctoral Research Excellence Scholarship in Imaging Science. Computing resources were provided by the University of British Columbia Cloud Innovation Centre, powered by Amazon Web Services https://cic.ubc.ca/. This research was also partly enabled by support provided by the Digital Research Alliance of Canada https://alliancecan.ca/, and the Research Computing Services group at the University of Calgary https://rcs.ucalgary.ca/RCS_Home_Page.